OpenStudy (anonymous):

In an experiment, a fair cubical dice was rolled repeatedly until a six resulted, and the number of rolls was recorded. The experiment was conducted 60 times. Find the expected frequency for two rolls correct to one decimal place.

OpenStudy (anonymous):

Is this question actually solvable?

OpenStudy (anonymous):

i don't know, i am bad probability hmm dalvoron can help you

OpenStudy (anonymous):

..at* probability

OpenStudy (anonymous):

Ok. I have an answer if it helps, but I just can't see how they got it... It's 8.3

OpenStudy (king):

probability is always lesser than 1 or equal to 1

OpenStudy (anonymous):

Yes, but frequency isn't...

OpenStudy (king):

oh i didn't see the question....sry

OpenStudy (anonymous):

"...and the number of rolls was recorded..." I am not able to get this! :/

OpenStudy (anonymous):

It just means that 60 rolls were recorded, as the second sentence says...

OpenStudy (anonymous):

I guess the expected frequency is the probability of a 6 occurring second (and only second), times the number of rolls. \[P(1,2,3,4,5)*P(6)*60\]\[=\frac{5}{6}\times\frac{1}{6}*60=8.3\]

OpenStudy (anonymous):

This makes a lot of sense... I thought something like that... the first roll was 1/6 and I didn't know that you could use a tree diagram to solve this... and then times it by number of times rolled... Yes it makes sense :) Thanks

OpenStudy (anonymous):

That probability comes from the situation where the first roll gets a not-6 (i.e. 1, 2, 3, 4, or 5), and the second roll gets a 6.

OpenStudy (anonymous):

wow Dalvoron

OpenStudy (anonymous):

Hugs for Ishaan!

OpenStudy (king):

awesome Dalvoron!!

OpenStudy (anonymous):

Thanks for solving :) the question was phrased in a strange way.

OpenStudy (anonymous):

thanks so much, this has helped me with my assignment.